# but Conditions Apply

### v1

@@ A.5.2 Implementation Details
Benchmarks.
Table A.3 presents the benchmarks used in our study, categorized into Attribute, Object, Position, Counting, Complex Relation, Action, Size, and others. The benchmarks include Vismin Awal et al. (2024), EQBench Wang et al. (2023), MMVP Tong et al. (2024), CLEVR Johnson et al. (2017), Whatsup Kamath et al. (2023), Spec Peng et al. (2024), ARO Yuksekgonul et al. (2023), Sugarcrepe Hsieh et al. (2023), COLA Ray et al. (2023), and Winoground Thrush et al. (2022).
Table A.3: Categorization of compositional benchmarks. For EQBench and Vismin, an official subset is used.
@@
$\mathbb{I}\Bigg{[}\begin{aligned} &\text{score}(\texttt{text1},\texttt{image1})>\text{score}(\texttt{text2},\texttt{image1})\\ &\hskip 79.66771pt\textbf{AND}\\ &\text{score}(\texttt{text2},\texttt{image2})>\text{score}(\texttt{text1},\texttt{image2})\end{aligned}\Bigg{]}$ (A.10)
@@
A.5.3 Details of Self-Bench
@@
Figure A.2 displays an example of the filtering interface where the annotator can label the images. Since there are ambiguous examples, such as an image with only half of an object, we provided three options: "Good", "Ambiguous", and "Wrong". If any one of the three annotators labels the example as ambiguous or wrong, the image is filtered out later. Examples of filtered images are provided in Figure A.3.
@@
Figure A.2: Example of a filtering userspace. The annotator should select one among three options. Before they annotate, they get the instructions for each category.
@@
Figure A.3: Images filtered out by annotators. Annotators removed images that did not meet the criteria specified in the category instructions.
A.6 Style alignment
Figure A.4: Image generations of SD3-m with style alignment. Top: CLEVR dataset; Bottom: Whatsup-A dataset.
The generations are shown in Fig. A.4. The results of these experiments are summarised in the table below. Overall, we find that style alignment through textual inversion may not be an effective way to mitigate the domain gap.
Table A.4: Diffusion classifier accuracy before and after style alignment via textual inversion.
Dataset Before Alignment (%) After Alignment (%)
WhatsApp-A 26.5 29.3
CLEVR-ColorBinding 59.1 57.1
A.7 Results on a distilled SD3.5 and FLUX model
This experiment tests whether aggressive distillation, while beneficial for generation quality and efficiency, compromises the discriminative performance of the model—or not. The results suggest that it does: the distilled model underperforms significantly compared to its base version. We present the results in Table A.5. Additionally, FLUX Yang et al. (2024) has been introduced with impressive generative quality. We include FLUX in our analysis as an exploratory case, focusing on Position and Counting tasks in real-world datasets—categories where diffusion models typically perform particularly well and poorly, respectively. As shown in Table A.6, we again observe a significant drop in discriminative performance for the distilled model.
Table A.5: SD3 / SD3.5 Large turbo (distilled version) accuracies on Self-Bench tasks.
Geneval Version Color Attr Position Counting Colors Single Two
1.5 0.17 / 0.15 0.67 / 0.63 0.33 / 0.30 0.53 / 0.72 0.48 / 0.74 0.19 / 0.49
2 0.31 / 0.34 0.47 / 0.48 0.41 / 0.42 0.65 / 0.71 0.51 / 0.78 0.16 / 0.57
3-m 0.49 / 0.72 0.40 / 0.68 0.43 / 0.57 0.83 / 0.84 0.73 / 0.88 0.43 / 0.75
Table A.6: Performance of FLUX (distilled version) in Position and Counting.
Method WhatsUP A WhatsUP B COCO one COCO two GQA one GQA two SPEC relative EQbench Vismin MMVP spatial MMVP orientation MMVP perspective CLEVR Vismin EQbench MMVP
SD1.5 0.28 0.27 0.48 0.52 0.49 0.47 0.30 0.15 0.19 0.27 0.07 0.53 0.49 0.33 0.25 0.33
SD2.0 0.27 0.28 0.42 0.47 0.49 0.53 0.29 0.15 0.13 0.33 0.13 0.6 0.51 0.13 0.15 0.13
SD3-m 0.28 0.37 0.54 0.55 0.54 0.56 0.43 0.05 0.44 0.47 0.33 0.47 0.59 0.13 0.05 0.13
FLUX 0.28 0.29 0.47 0.55 0.52 0.50 0.35 0.0 0.24 0.27 0.2 0.33 0.57 0.17 0.05 0.13
A.8 Additional qualitative results
We illustrate a qualitative example of SD2.0 and SD3-m generation results from different timesteps on Whatsup-A dataset in Figure A.5.
Figure A.5: Image generation results on Whatsup-A using SD2.0 and SD3-m models.
A.9 On universality of timestep weighting
We present the results in Table A.7. Overall, we find that uniform and exponentially weighted models perform quite similarly for 1.52 diffusion models, although the gap is often quite large between the two. While for SD-3m model, uniform weighting is almost always better.
Table A.7: Mean accuracies by version and task. Bold indicates the larger of the uniform or exponentially–weighted scores.
Version Task Uniform Exp. Weighted
1.5 COCO QA 0.44 0.48
VG QA 0.44 0.49
CLEVR Binding–Color 0.67 0.70
CLEVR Spatial 0.50 0.48
Spec Count 0.20 0.12
Sugar Attributes 0.70 0.61
Sugar Objects 0.85 0.78
WhatsUp A 0.27 0.27
WhatsUp B 0.26 0.28
2 COCO QA 0.42 0.47
VG QA 0.48 0.49
CLEVR Binding–Color 0.82 0.74
CLEVR Spatial 0.49 0.50
Spec Count 0.23 0.12
Sugar Attributes 0.76 0.64
Sugar Objects 0.85 0.81
WhatsUp A 0.26 0.30
WhatsUp B 0.26 0.23
3-m COCO QA 0.56 0.55
VG QA 0.54 0.52
CLEVR Binding–Color 0.60 0.75
CLEVR Spatial 0.62 0.51
Spec Count 0.19 0.11
Sugar Attributes 0.72 0.61
Sugar Objects 0.74 0.71
WhatsUp A 0.28 0.27
WhatsUp B 0.40 0.29
A.10 Upperbounding Self-Bench performance with timestep weighting
Figure A.6: Self-Bench: Timestep reweighting helps address the cross-domain problem. We show the performance of each SD model on cross-domain data (e.g., for the SD1.5 model, the bars depict its performance on SD2.0 and SD3-m generations, averaged). All models benefit from timestep reweighting, with SD3-m benefiting the most.
Our upper bound analysis in Self-Bench reveals that optimizing timestep weights improves performance across most tasks and models (Figure A.6). The improvements are particularly dramatic for SD3-m: in spatial tasks, accuracy increases from 33% to 95%, and in the two-objects task from 41% to 73%. While SD1.5 and SD2.0 also benefit from reweighting, showing improvements of 1-10% across tasks, the gains are more modest compared to SD3-m. This suggests that SD3-m’s lower baseline performance is not due to fundamental model limitations, but rather suboptimal weighting of timestep information.
A.11 Later timesteps
We explore whether further gains could be achieved with SD3 based on our prior analysis. Motivated by findings from timestep weighting, we tested a simplified variant where only mid-to-late timesteps were used for discrimination—specifically in cases where either single-timestep accuracy or learned weights emphasized later timesteps. In conclusion, however, these simplified tests did not yield meaningful improvements. The results are presented in Table A.8.
Table A.8: Results using later timesteps (sampling from $t\sim[0.5,1]$ for classification.
SD3-m 0.87 0.57 0.30 0.63
SD3-m (supervised) 0.91 0.72 0.42 0.98
Later Timesteps 0.90 0.66 0.31 0.59
A.12 CLIP scores and domain gaps
As an extension of Figure 9 in Section 5.4 of the main paper, Table A.9 shows CLIP embedding distances between real-world datasets and Self-Bench generations, and corresponding accuracy gains from timestep weighting. We can see the positive correlation in SD3 but not in SD1 and SD2.
Table A.9: Timestep Weighting and Domain Gap. CLIP embedding distances between real-world datasets and Self-Bench generations, and corresponding accuracy gains from timestep weighting.
COCO QA 3.494 5% 3.133 5% 3.576 +4%
VQ QA 3.666 5% 3.584 1% 3.918 +5%
SPEC Count 2.646 2% 2.598 0% 3.191 +5%
WhatsUp B 4.254 0% 4.016 0% 4.047 +4%
WhatsUp A 4.656 -1% 4.344 4% 4.600 +12%
CLEVR Binding 5.64 9% 4.348 8% 4.926 +35%
CLEVR Spatial 5.57 -3% 4.63 2% 5.023 +16%
  • $\ddagger$
    Position  $\dagger$ Counting  $\star$ Attribute
A.13 Additional results
Qualitative examples. Figure A.7 illustrates the domain differences between existing compositional benchmarks and SD3-m generated results, using prompts from Self-Bench, SPEC Peng et al. (2024), and WhatsUP Kamath et al. (2023). Figure A.8 presents additional examples from Self-Bench. Figure A.9 highlights failure cases of the Diffusion Classifier, even in in-domain scenarios.
Compositional Benchmarks.
Table A.10A.11A.12A.13A.14A.15A.16A.17A.18 and A.19 report quantitative results for Winoground Thrush et al. (2022), COLA Ray et al. (2023), EQBench Wang et al. (2023), VisMin Awal et al. (2024), MMVP Tong et al. (2024), ARO Yuksekgonul et al. (2023), CLEVR Johnson et al. (2017), SugarCrepe Hsieh et al. (2023), WhatsUP Kamath et al. (2023), and SPEC Peng et al. (2024), respectively. Figure A.10 provides an overview of results across all prior benchmarks considered in our study.
Self-Bench. Table A.20 summarizes performance on Self-Bench. Given its high accuracy, one might suspect data bias in Self-Bench or selection bias in the model. To address this, we report both macro and micro accuracy in Figure A.11, which also provides a comprehensive overview of performance on Self-Bench. Figure A.12 shows performance degradation in cross-domain settings. Table A.21, Table A.22, and Table A.23 present fine-grained performance across color, position, and counting tasks, respectively.
Compositional Benchmarks vs Self-Bench. Figure A.13 compares results from existing compositional benchmarks with those from Self-Bench, revealing a clear domain shift across all tasks (see gray background).
Figure A.7: Domain differences between existing compositional benchmarks and SD3-m generated results. (top) Examples from existing compositional benchmarks: SPEC Peng et al. (2024) and WhatsUP A Kamath et al. (2023). (bottom left) Self-Bench examples. (bottom middle) Generated images using prompts from SPEC count tasks. (bottom right) Generated images using prompts from WhatsUPA tasks.
@@ A.13 Additional results
Figure A.8: More examples from Self-Bench. Representative samples illustrating the range of tasks and model outputs used in our benchmark.
@@
Figure A.9: Self-Bench in-Domain prediction failure cases. Examples where the model fails despite being evaluated in an in-domain setting.
@@
Figure A.10: Overview of performance on compositional benchmarks beyond the four main categories. Summarizes model accuracy on additional compositional tasks, highlighting generalization beyond the core evaluation categories. The red dotted horizontal line represents the random chance level.
Table A.10: Image-to-text retrieval results on Winoground Thrush et al. (2022).
@@
Table A.11: Image-to-Text retrieval accuracy on COLA Ray et al. (2023).
@@
Table A.12: Image-to-text retrieval accuracy on EQbench Wang et al. (2023).
@@
Table A.13: Image-to-text retrieval accuracy on VisminAwal et al. (2024).
@@
Table A.14: Image-to-Text retrieval accuracy on MMVP-VLM Tong et al. (2024).
@@
Table A.15: Image-to-Text retrieval accuracy on ARO Yuksekgonul et al. (2023).
@@
Table A.16: Image-to-text retrieval accuracy on CLEVR Johnson et al. (2017).
@@
Table A.17: Image-to-Text retrieval accuracy on SugarCrepe Hsieh et al. (2023).
@@
Table A.18: Image-to-Text retrieval accuracy on WhatsUP Kamath et al. (2023).
@@
Table A.19: Image-to-text retrieval accuracy on SPEC Peng et al. (2024).
@@
Figure A.11: Micro accuracy (Top) and Macro accuracy (Bottom) on self-Bench. We evaluate CLIP and SD models in our Self-Bench. The X-axis represents the model used for generating the dataset, with the number of images for each dataset indicated below. The results clearly show that models perform well only when evaluated in-domain, not cross-domain. The red dotted horizontal line represents the random chance level.
Figure A.12: Self-Bench cross-domain drop rate. WE show the performance drop when evaluating models in cross-domain settings.
Table A.20: Complete image-to-text retrieval accuracy results on Self-Bench across all tasks. Bold entries indicate in-domain evaluations for the model.
@@
Table A.21: Fine-grained accuracy on Self-Bench Colors.
@@
Table A.22: Fine-grained accuracy on Self-Bench Position
@@
Table A.23: Fine-grained accuracy on Self-Bench Counting
@@
Figure A.13: Additional results on compositional benchmarks and Self-Bench. The detailed breakdown of our analysis within four categories: (a) Object, (b) Attribute, (c) Position, and (d) Counting. Results on a white background correspond to performance across ten existing compositional benchmarks (33 sub-tasks), while those on a gray background represent results on Self-Bench.The red dotted horizontal line represents the random chance level.

### v3

@@ C.2 Implementation Details
Benchmarks. Table C.1 presents the benchmarks used in our study, categorized into Attribute, Object, Position, Counting, Complex Relation, Action, Size, and others. The benchmarks include Vismin Awal et al. (2024), EQBench Wang et al. (2023), MMVP Tong et al. (2024), CLEVR Johnson et al. (2017), Whatsup Kamath et al. (2023), Spec Peng et al. (2024), ARO Yuksekgonul et al. (2023), Sugarcrepe Hsieh et al. (2023), COLA Ray et al. (2023), and Winoground Thrush et al. (2022).
Table C.1: Categorization of compositional benchmarks. For EQBench and Vismin, an official subset is used.
@@
$\mathbb{I}\Bigg[\begin{aligned} &\text{score}(\texttt{text1},\texttt{image1})>\text{score}(\texttt{text2},\texttt{image1})\\ &\hskip 79.66771pt\textbf{AND}\\ &\text{score}(\texttt{text2},\texttt{image2})>\text{score}(\texttt{text1},\texttt{image2})\end{aligned}\Bigg]$ (C.1)
@@
Appendix D Details of Self-Bench
D.1 Design and filtering
@@
Figure D.1 displays an example of the filtering interface where the annotator can label the images. Since there are ambiguous examples, such as an image with only half of an object, we provided three options: "Good", "Ambiguous", and "Wrong". If any one of the three annotators labels the example as ambiguous or wrong, the image is filtered out later. Examples of filtered images are provided in Figure D.2.
@@
Figure D.1: Example of a filtering userspace. The annotator should select one among three options. Before they annotate, they get the instructions for each category.
@@
Figure D.2: Images filtered out by annotators. Annotators removed images that did not meet the criteria specified in the category instructions.
D.2 Qualitative examples
Figure D.3 illustrates the domain differences between existing compositional benchmarks and SD3-m generated results, using prompts from Self-Bench, SPEC Peng et al. (2024), and WhatsUP Kamath et al. (2023). Figure D.4 presents additional examples from Self-Bench. Figure D.5 highlights failure cases of the Diffusion Classifier, even in in-domain scenarios.
Figure D.3: Domain differences between existing compositional benchmarks and SD3-m generated results. (top) Examples from existing compositional benchmarks: SPEC Peng et al. (2024) and WhatsUP A Kamath et al. (2023). (bottom left) Self-Bench examples. (bottom middle) Generated images using prompts from SPEC count tasks. (bottom right) Generated images using prompts from WhatsUPA tasks.
@@ D.2 Qualitative examples
Figure D.4: More examples from Self-Bench. Representative samples illustrating the range of tasks and model outputs used in our benchmark.
@@
Figure D.5: Self-Bench in-Domain prediction failure cases. Examples where the model fails despite being evaluated in an in-domain setting.
Appendix E Additional results
E.1 Full results in compositional benchmarks
Table E.1E.2E.3E.4E.5E.6E.7E.8E.9 and E.10 report quantitative results for Winoground Thrush et al. (2022), COLA Ray et al. (2023), EQBench Wang et al. (2023), VisMin Awal et al. (2024), MMVP Tong et al. (2024), ARO Yuksekgonul et al. (2023), CLEVR Johnson et al. (2017), SugarCrepe Hsieh et al. (2023), WhatsUP Kamath et al. (2023), and SPEC Peng et al. (2024), respectively. Figure E.1 provides an overview of results across all prior benchmarks considered in our study.
Self-Bench. Table E.11 summarizes performance on Self-Bench. Given its high accuracy, one might suspect data bias in Self-Bench or selection bias in the model. To address this, we report both macro and micro accuracy in Figure E.2, which also provides a comprehensive overview of performance on Self-Bench. Figure E.3 shows performance degradation in cross-domain settings. Table E.12, Table E.13, and Table E.14 present fine-grained performance across color, position, and counting tasks, respectively.
Compositional Benchmarks vs Self-Bench. Figure E.4 compares results from existing compositional benchmarks with those from Self-Bench, revealing a clear domain shift across all tasks (see gray background).
@@ E.1 Full results in compositional benchmarks
Figure E.1: Overview of performance on compositional benchmarks beyond the four main categories. Summarizes model accuracy on additional compositional tasks, highlighting generalization beyond the core evaluation categories. The red dotted horizontal line represents the random chance level.
Table E.1: Image-to-text retrieval results on Winoground Thrush et al. (2022).
@@
Table E.2: Image-to-Text retrieval accuracy on COLA Ray et al. (2023).
@@
Table E.3: Image-to-text retrieval accuracy on EQbench Wang et al. (2023).
@@
Table E.4: Image-to-text retrieval accuracy on VisminAwal et al. (2024).
@@
Table E.5: Image-to-Text retrieval accuracy on MMVP-VLM Tong et al. (2024).
@@
Table E.6: Image-to-Text retrieval accuracy on ARO Yuksekgonul et al. (2023).
@@
Table E.7: Image-to-text retrieval accuracy on CLEVR Johnson et al. (2017).
@@
Table E.8: Image-to-Text retrieval accuracy on SugarCrepe Hsieh et al. (2023).
@@
Table E.9: Image-to-Text retrieval accuracy on WhatsUP Kamath et al. (2023).
@@
Table E.10: Image-to-text retrieval accuracy on SPEC Peng et al. (2024).
@@
Figure E.2: Micro accuracy (Top) and Macro accuracy (Bottom) on self-Bench. We evaluate CLIP and SD models in our Self-Bench. The X-axis represents the model used for generating the dataset, with the number of images for each dataset indicated below. The results clearly show that models perform well only when evaluated in-domain, not cross-domain. The red dotted horizontal line represents the random chance level.
Figure E.3: Self-Bench cross-domain drop rate. WE show the performance drop when evaluating models in cross-domain settings.
Table E.11: Complete image-to-text retrieval accuracy results on Self-Bench across all tasks. Bold entries indicate in-domain evaluations for the model.
@@
Table E.12: Fine-grained accuracy on Self-Bench Colors.
@@
Table E.13: Fine-grained accuracy on Self-Bench Position
@@
Table E.14: Fine-grained accuracy on Self-Bench Counting
@@
Figure E.4: Additional results on compositional benchmarks and Self-Bench. The detailed breakdown of our analysis within four categories: (a) Object, (b) Attribute, (c) Position, and (d) Counting. Results on a white background correspond to performance across ten existing compositional benchmarks (33 sub-tasks), while those on a gray background represent results on Self-Bench.The red dotted horizontal line represents the random chance level.
E.2 Results on a distilled SD3.5 and FLUX model
This experiment tests whether aggressive distillation, while beneficial for generation quality and efficiency, compromises the discriminative performance of the model—or not. The results suggest that it does: the distilled model underperforms significantly compared to its base version. We present the results in Table E.15. Additionally, FLUX Yang et al. (2024) has been introduced with impressive generative quality. We include FLUX in our analysis as an exploratory case, focusing on Position and Counting tasks in real-world datasets—categories where diffusion models typically perform particularly well and poorly, respectively. As shown in Table E.16 and Table E.17, we again observe a significant drop in discriminative performance for the distilled model.
Table E.15: SD3 Large turbo (distilled) / SD3-m accuracies on Self-Bench tasks.
Geneval Version Color Attr Position Counting Colors Single Two
1.5 0.17 / 0.15 0.67 / 0.63 0.33 / 0.30 0.53 / 0.72 0.48 / 0.74 0.19 / 0.49
2 0.31 / 0.34 0.47 / 0.48 0.41 / 0.42 0.65 / 0.71 0.51 / 0.78 0.16 / 0.57
3-m 0.49 / 0.72 0.40 / 0.68 0.43 / 0.57 0.83 / 0.84 0.73 / 0.88 0.43 / 0.75
Table E.16: Performance of FLUX in Position.
Method WhatsUP A WhatsUP B COCO one COCO two GQA one GQA two SPEC relative EQbench Vismin MMVP spatial MMVP orientation MMVP perspective CLEVR SPEC absolute
SD1.5 0.28 0.27 0.48 0.52 0.49 0.47 0.30 0.15 0.19 0.27 0.07 0.53 0.49 0.15
SD2.0 0.27 0.28 0.42 0.47 0.49 0.53 0.29 0.15 0.13 0.33 0.13 0.6 0.51 0.12
SD3-m 0.28 0.37 0.54 0.55 0.54 0.56 0.43 0.05 0.44 0.47 0.33 0.47 0.59 0.24
FLUX 0.28 0.29 0.47 0.55 0.52 0.50 0.35 0.0 0.24 0.27 0.2 0.33 0.57 0.17
Table E.17: Performance of FLUX in Counting.
Method Vismin EQbench MMVP SPEC
SD1.5 0.33 0.25 0.33 0.2
SD2.0 0.13 0.15 0.13 0.23
SD3-m 0.13 0.05 0.13 0.18
FLUX 0.17 0.05 0.13 0.15
E.3 Style alignment
Figure E.5: Image generations of SD3-m with style alignment. Top: CLEVR dataset; Bottom: Whatsup-A dataset.
The generations are shown in Fig. E.5. The results of these experiments are summarised in the table below. Overall, we find that style alignment through textual inversion may not be an effective way to mitigate the domain gap.
Table E.18: Diffusion classifier accuracy before and after style alignment via textual inversion.
Dataset Before Alignment (%) After Alignment (%)
WhatsApp-A 26.5 29.3
CLEVR-ColorBinding 59.1 57.1
E.4 Additional qualitative results
We illustrate a qualitative example of SD2.0 and SD3-m generation results from different timesteps on Whatsup-A dataset in Figure E.6.
Figure E.6: Image generation results on Whatsup-A using SD2.0 and SD3-m models.
E.5 Upperbounding Self-Bench performance with timestep weighting
Figure E.7: Self-Bench: Timestep reweighting helps address the cross-domain problem. We show the performance of each SD model on cross-domain data (e.g., for the SD1.5 model, the bars depict its performance on SD2.0 and SD3-m generations, averaged). All models benefit from timestep reweighting, with SD3-m benefiting the most.
Our upper bound analysis in Self-Bench reveals that optimizing timestep weights improves performance across most tasks and models (Figure E.7). The improvements are particularly dramatic for SD3-m: in spatial tasks, accuracy increases from 33% to 95%, and in the two-objects task from 41% to 73%. While SD1.5 and SD2.0 also benefit from reweighting, showing improvements of 1-10% across tasks, the gains are more modest compared to SD3-m. This suggests that SD3-m’s lower baseline performance is not due to fundamental model limitations, but rather suboptimal weighting of timestep information.
E.6 CLIP scores and domain gaps
As an extension of Figure 9 in Section 5.4 of the main paper, Table E.19 shows CLIP embedding distances between real-world datasets and Self-Bench generations, and corresponding accuracy gains from timestep weighting. We can see the positive correlation in SD3 but not in SD1 and SD2.
Table E.19: Timestep Weighting and Domain Gap. CLIP embedding distances between real-world datasets and Self-Bench generations, and corresponding accuracy gains from timestep weighting.
COCO QA 3.494 5% 3.133 5% 3.576 +4%
VQ QA 3.666 5% 3.584 1% 3.918 +5%
SPEC Count 2.646 2% 2.598 0% 3.191 +5%
WhatsUp B 4.254 0% 4.016 0% 4.047 +4%
WhatsUp A 4.656 -1% 4.344 4% 4.600 +12%
CLEVR Binding 5.64 9% 4.348 8% 4.926 +35%
CLEVR Spatial 5.57 -3% 4.63 2% 5.023 +16%
  • $\ddagger$
    Position   $\dagger$ Counting   $\star$ Attribute
E.7 SigLIP results
We also report SigLIP and SigLIP2 (ViT‑SO400M‑14 and ViT‑L‑16‑256) as additional discriminative baselines in Table E.20. While SigLIP sometimes exceeds CLIP (e.g., SugarCrepe‑Attribute), the conclusion remains the same: classifiers remain best in Position, and CLIP/SigLIP dominate Object/Counting.
Table E.20: SigLIP baselines. We report the best CLIP variant, SigLIP, SigLIP2, and the best diffusion classifier per task.
Benchmark Best Diffusion Best CLIP SigLIP SigLIP2
Self-Bench (1.5) 0.88 0.79 0.80 0.81
Self-Bench (2.0) 0.94 0.84 0.86 0.89
Self-Bench (3-m) 0.95 0.80 0.79 0.82
COCO QA two 0.55 0.53 0.50 0.49
VQ QA two 0.56 0.55 0.49 0.51
SPEC Count 0.23 0.47 0.40 0.41
WhatsUP A 0.28 0.34 0.28 0.30
WhatsUP B 0.37 0.31 0.28 0.28
CLEVR Colors Bind. 0.81 0.53 0.50 0.51
CLEVR Spatial 0.59 0.54 0.50 0.50
SugarCrepe attrib. 0.75 0.75 0.78 0.77
SugarCrepe object 0.87 0.92 0.91 0.92